OpenStudy (anonymous):

simply tan^2x/secx Please someone answer...asap with all steps

OpenStudy (anonymous):

\[\tan^2(x)/\sec(x)\] \[=\frac{ \sin^2(x) }{ \cos^2(x) } * \frac{ 1 }{ \cos(x) }\] this simplifies down to \[= \frac{ \sin^2(x) }{ \cos(x) } = \sin(x) *\tan (x) = \sec(x)-\cos(x)\]

OpenStudy (cggurumanjunath):